a. The price of a cup of coffee was $0.29 in 1970 and $0.99 in 1993. The CPI was 38.8 in 1970 and 144.0 in 1993. Calculate the 1993 price of a cup of coffee in 1970 dollars.
b. For Chinese consumers, the income elasticity of demand for fruit juice is 1.1. If the economy enters a recession next year and consumer income declines by 2.5%, what is the expected change in the quantity of fruit juice demanded next year?
1
Expert's answer
2019-11-11T15:44:38-0500

Solution:

a) Calculate the 1993 price of a cup of coffee in 1970 dollars


"p= \frac{p1970\times CPI1993}{p1993\times CPI 1970}"

"p=\frac{0.29\times144}{0.99\times33.8}=1.25"

Answer:1.25

b) The expected change in the quantity of fruit juice demanded next year is:


"\varDelta Q=E\times \varDelta I"

"\varDelta Q=1.1\times(-2.5)=-2.75"

Answer: Will decrease by 2.75%